# Langevin equations and a geometric integration scheme for the overdamped limit of homogeneous rotational Brownian motion

### v2

@@ Abstract
The translational motion of anisotropic or self-propelled colloidal particles is closely linked with the particle’s orientation and its rotational Brownian motion. In the overdamped limit, the stochastic evolution of the orientation vector follows a diffusion process on the unit sphere and is characterised by an orientation-dependent (“multiplicative”) noise. As a consequence, the corresponding Langevin equation attains different forms depending on whether Itō’s or Stratonovich’s stochastic calculus is used. We clarify that both forms are equivalent and derive them from a geometric construction of Brownian motion on the unit sphere, based on infinitesimal random rotations. Our approach suggests further a geometric integration scheme for rotational Brownian motion, which preserves the normalisation constraint of the orientation vector exactly. We show that a simple implementation of the scheme, based on Gaussian random rotations, converges weakly at order 1 of the integration time step, and we outline an advanced variant of the scheme that is weakly exact for an arbitrarily large time step. Due to a favourable prefactor of the discretisation error, already the Gaussian scheme allows for integration time steps that are one order of magnitude larger compared to a commonly used, ad hoc algorithm for rotational Brownian dynamics simulations. The discussion is restricted to time-homogeneous rotational Brownian motion (i.e., a single rotational diffusion constant), which is relevant for axisymmetric particles and also chemically anisotropic spheres such as self-propelled Janus particles.
@@ I Introduction
The past two decades have seen considerable advances in the synthesis and characterisation of anisotropic colloidal particles, ranging from complex shapes, e.g., ellipsoids [1, 2], clusters [3], and propellers [4], to anisotropic chemical surface coatings. An important example for the latter are so-called Janus spheres [5, 6, 7, 8, 9, 10], which exhibit autonomous self-propulsion under suitable conditions and often serve as colloidal models for microswimmers. The use of such particles in experimental model systems has opened new avenues and has boosted research in the field of active matter (see, e.g., Refs. [11, 12, 13]). These successes have stimulated a large number of theoretical investigations, which have brought forth elaborate descriptions of the self-propelled motion of a single active colloid (see, e.g., Refs. [14, 15, 16, 17, 18, 19]). Depending on the question at hand, hydrodynamic effects due to solvent flows can be crucial [20, 21, 22, 23] or may be negligible. In the latter case, the active Brownian particle (ABP) model is widely used to describe the motion of active colloids [24, 17, 25]; it combines the Brownian motion of a colloidal Janus sphere with a constant propulsion velocity which randomly changes its direction. The propulsion direction is described by a unit vector $\boldsymbol{u}$ which itself performs a Brownian motion on the unit sphere. Thereby, the stochastic motion of $\boldsymbol{u}$ may evolve freely [5, 26, 27] or under the influence of gravity [28, 29, 30, 31, 32, 33] and it becomes particularly challenging in the case of shape anisotropy [34, *Doi:2005, 36, 37, 38, 39, 40, 41, *Leitmann:2017, 43]. The dynamics of active colloids becomes complex in the presence of rotation–translation coupling, e.g., due to short-time hydrodynamic friction [34, *Doi:2005, 36, 37] or due to confinement by a harmonic potential well [44] or a substrate potential [27, 45]; in such situations, the description of the rotational dynamics is crucial for predictions on the translational motion.
@@
Models for the rotational motion of molecules have traditionally been studied in the context of dynamic light scattering and dielectric spectroscopy [46, 47]. These models are typically based on the Fokker–Planck equation for orientational diffusion [48, 49, 50], which allows one to calculate correlation functions and, more generally, the statistics of the trajectories [40, 41, *Leitmann:2017, 43]. Alternatively, some form of Langevin equation is employed, which provides a description on the level of stochastic trajectories and which forms the basis of stochastic simulations. The underdamped Langevin equation is used when inertia effects are relevant [46] and, mathematically, there are no conceptual questions with this equation apart from being demanding to work with it analytically. In the context of colloidal motion, however, one would like to model the stochastic dynamics as (completely) overdamped, which requires a suitable Wiener process for the particle orientation. The corresponding Langevin equation contains a noise term that has a “multiplicative” structure [46, 34, *Doi:2005, 36, 37, 29, 31, 25, 41, *Leitmann:2017, 43, 27, 45] [see below, Eqs. 7 and 9], which brings up questions about the underlying stochastic calculus. In particular, in its Itō form, the overdamped Langevin equation [Eq. 9] contains a drift-like term, which may appear counter-intuitive at first sight. Another example for this mathematically well-understood situation [51, 52] is given by translational diffusion near a wall characterised by a position-dependent diffusivity, which has lead to a re-arising apparent controversy over decades [53, 54, 55, 56, 57, 58, 59, 60, 61, 62]. The issue of multiplicative noise in rotational diffusion is entirely circumvented by modelling the particle orientation as confined to a plane, which allows one to parametrise the two-dimensional rotations by a single angle. Whereas this choice is widely adopted in models of active matter, it often appears to be a choice of convenience for simplicity rather than being physically well justified. Notably, the linear transport behavior can sensitively depend on the dimensionality of rotational diffusion [45].
@@
In this note, we aim to clarify some mathematical aspects of rotational Brownian motion, in particular, the different forms of the overdamped Langevin equation and how they depend on the adopted definition of the stochastic integral (Itō or Stratonovich). The equations are derived from a geometric construction of Brownian motion on the unit sphere, given as a sequence of infinitesimal random rotations. The construction suggests a geometric integration scheme for rotational Brownian motion, which (i) exactly preserves the normalisation constraint of the orientation vector and (ii) can be extended such that it is weakly exact for arbitrarily large time steps. In order to focus on the essence of the problem, we shall mainly discuss chemically anisotropic spheres (e.g., Janus spheres), thereby avoiding additional difficulties that arise from the incessant transformation of the diffusivity tensor due to the rotational motion. We then show that the results apply also to axisymmetric particles, however, their derivation requires some modifications which are mostly of technical nature. We refer to a Brownian motion with constant coefficients as homogeneous in time.
@@
Finally, we note that the literature describes a number of numerical schemes for Brownian motion of rigid bodies, which are often justified only heuristically; this includes schemes using finite rotations [63, 64], adding a constraint force [65, 66, 67, 68], discretising the Stratonovich–Langevin equation [34, 69], or employing a symplectic splitting scheme of the underdamped Langevin dynamics [70]. Also the algorithm presented below (Section IV.1) has already been used in previous work of both authors [41, *Leitmann:2017, 27, 45], without further explanation.
@@ II.1 Langevin equation of rotational Brownian motion with inertia
If the particle is immersed in a fluid medium, it experiences a torque $-\boldsymbol{\zeta}_{R}(t)\,\boldsymbol{\omega}(t)$ due to viscous (Stokes) friction. In general, the rotational friction coefficient $\boldsymbol{\zeta}_{R}(t)$ is a tensor which is constant in a body-fixed frame [71], but is time-dependent in the space-fixed frame used here. Following Langevin’s and Ornstein’s approach to Brownian motion, the incessant collisions of the particle with solvent molecules are further subsumed in a random torque $\boldsymbol{\xi}(t)$ acting on the particle. Taking the collisions to be uncorrelated in a first approximation suggests to model $\boldsymbol{\xi}(t)$ as a Gaussian white noise process, which is independent of $\boldsymbol{\omega}(s)$ for $t>s$, has mean zero, and its covariance matrix is determined by the fluctuation–dissipation relation:
@@
This is a stochastic differential equation in $\boldsymbol{\omega}(t)$ and since the white noise $\boldsymbol{\xi}(t)$ enters “additively” (i.e., it is not multiplied by a $\boldsymbol{\omega}(t)$-dependent function), there is no ambiguity about the choice of the stochastic calculus: any (consistent) definition of the stochastic integral yields the same solution [52].
In Section V, we extend the results to axisymmetric particles, for which $\boldsymbol{u}(t)$ is naturally chosen to align with the symmetry axis of the particle. This axis fully determines the particle’s orientation such that the friction tensor $\boldsymbol{\zeta}_{R}(t)$ and moment of inertia $\boldsymbol{I}(t)$ are functions of $\boldsymbol{u}(t)$. For particles of arbitrary shape, the orientation may be tracked by following the motion of a body-fixed trihedron [34, *Doi:2005, 36], an elaboration on this generalisation is left for future research.
II.2 Heuristic derivation of the overdamped limit
[…]
This simple reasoning is possible since Eq. 4 is of the Ornstein–Uhlenbeck type, in particular, since it is linear in $\boldsymbol{\omega}(t)$ and the noise $\boldsymbol{\xi}(t)$ enters additively. The same result would be obtained from the formal solution for $\boldsymbol{\omega}(t)$, given in integral form [52], and taking the limit $\tau_{1}\to 0$.
Equation 5 states that the angular velocity $\boldsymbol{\omega}(t)$ is statistically equivalent to a Gaussian white noise process with a deterministic bias $\boldsymbol{T}_{\text{ext}}(t)$. More precisely, we may take $\boldsymbol{\omega}(t)$ to be a Gaussian process which satisfies
$\langle\delta\boldsymbol{\omega}(t^{\prime})\otimes\delta\boldsymbol{\omega}(t)\rangle$ $=2D_{\text{R}}\boldsymbol{1}\delta(t^{\prime}-t)$ (6)
[…]
It is now natural to interpret the kinematic equation (2) for the orientation $\boldsymbol{u}(t)$ as a stochastic differential equation with a random angular velocity $\boldsymbol{\omega}(t)$. We repeat the equation here:
[…]
which is one form of the overdamped Langevin equation for rotational Brownian motion of a body-fixed axis $\boldsymbol{u}(t)$. However, the white noise $\boldsymbol{\omega}(t)$ multiplies the variable $\boldsymbol{u}(t)$ and the re-interpretation of Eq. 2 is yet ill-defined unless we specify the stochastic calculus  [^1] We emphasise that we have left open this question when solving Eq. 4., i.e., unless we give a meaning to the time integral $\int_{0}^{t}\boldsymbol{\omega}(s)\times\boldsymbol{u}(s)\,\mathrm{d}s$. This is a modelling decision. Recalling that $\boldsymbol{u}(t)$ is a unit vector, Eq. 7 describes a diffusion process on a constraining manifold, and the natural choice in this situation is to interpret the time integral as a (Fisk–)Stratonovich integral (see chap. 35 in ref. 79).
In the physics literature, the Stratonovich interpretation of Eq. 7 is often (tacitly and correctly) assumed. For the numerical integration, however, a direct application of the widespread Euler–Maruyama algorithm, which is a simple and robust algorithm for Itō SDEs, would break the conservation of the norm $|\boldsymbol{u}(t)|$. This could be fixed a posteriori by rescaling $\boldsymbol{u}(t)$ after every integration step [66], but such an approach appears conceptually unsatisfying  [^2] In fact, this debatable ad hoc algorithm is currently implemented in the molecular dynamics simulation software LAMMPS (since release from 29 September 2021), see the command “fix Brownian/sphere” [see https://docs.lammps.org/fix_brownian.html]. . Instead, one should use a suitable discretisation scheme for the Stratonovich integral, e.g., the stochastic Milstein algorithm; yet, the correct implementation of such schemes is more demanding [81, 68]. A conceptually less satisfying alternative is to mix different stochastic calculi: one may cast the Stratonovich form (7) into an equivalent form that invokes the Itō integral and which is thus suitable for the Euler–Maruyama scheme [82, 46, 52]; this form includes a noise-induced term appearing as a drift [see below, Eq. 9]. We note that all of these integration schemes preserve the normalisation of $\boldsymbol{u}(t)$ only asymptotically for a vanishing integration time step, $\Delta t\to 0$.
II.3 Two-dimensional case.
For motion in two dimensional space, which is widely considered in the theoretical literature on active particles, the issue of the multiplicative noise can be circumvented since the angular velocity has only a single component. Writing $\boldsymbol{\omega}=(0,0,\omega)$, the component $u_{z}$ of $\boldsymbol{u}=(u_{x},u_{y},u_{z})$ remains unchanged under the dynamics so that we can set $u_{z}=0$ and parametrise the remaining two components $(u_{x},u_{y})=(\cos(\varphi),\sin(\varphi))$ in terms of the angle $\varphi\in[-\pi,\pi)$. The evolution of $\varphi(t)$ follows from Eq. 2, which reads for the first vector component: $\dot{u}_{x}(t)=-\omega\,u_{y}(t).$ Combining with the ordinary chain rule, $\dot{u}_{x}(t)=-\sin(\varphi(t))\,\dot{\varphi}(t),$ yields:
$\dot{\varphi}(t)=\omega(t)\,.$ (8)
Passing to the overdamped limit, we replace $\omega(t)$ with the white noise given in Eq. 5, which enters additively. In fact, Eq. 8 is the simplest form of a stochastic differential equation, which has an unambiguous solution with any stochastic calculus. We conclude that $\varphi(t)$ is a Brownian motion on the one-dimensional torus $[-\pi,\pi)$, i.e., $\boldsymbol{u}(t)$ performs a Brownian motion on the unit circle in the $xy$-plane. The same result would be obtained when starting from Eq. 7 and observing that the classical (Leibniz) chain rule still applies for the Stratonovich calculus.
III Geometric derivation of the overdamped Langevin equation
For analytical and numerical work the use of Itō’s calculus has certain advantages and, to this end, Eq. 7 needs to be cast into its Itō form, which reads:
[…]

### v3

# Langevin equations and a geometric integration scheme for the overdamped limit of rotational Brownian motion of axisymmetric particles

@@ Abstract
The translational motion of anisotropic or self-propelled colloidal particles is closely linked with the particle’s orientation and its rotational Brownian motion. In the overdamped limit, the stochastic evolution of the orientation vector follows a diffusion process on the unit sphere and is characterized by an orientation-dependent (“multiplicative”) noise. As a consequence, the corresponding Langevin equation attains different forms depending on whether Itō’s or Stratonovich’s stochastic calculus is used. We clarify that both forms are equivalent and derive them in a top-down appraoch from a geometric construction of Brownian motion on the unit sphere, based on infinitesimal random rotations. Our approach suggests further a geometric integration scheme for rotational Brownian motion, which preserves the normalization constraint of the orientation vector exactly. We show that a simple implementation of the scheme, based on Gaussian random rotations, converges weakly at order 1 of the integration time step, and we outline an advanced variant of the scheme that is weakly exact for an arbitrarily large time step. Due to a favorable prefactor of the discretization error, already the Gaussian scheme allows for integration time steps that are one order of magnitude larger compared to a commonly used algorithm for rotational Brownian dynamics simulations based on projection on the constraining manifold. For torques originating from constant external fields, we prove by virtue of the Fokker–Planck equation that the constructed diffusion process satisfies detailed balance and converges to the correct equilibrium distribution. The analysis is restricted to time-homogeneous rotational Brownian motion (i.e., a single rotational diffusion constant), which is relevant for axisymmetric particles and also chemically anisotropic spheres, such as self-propelled Janus particles.
@@ I Introduction
The past two decades have seen considerable advances in the synthesis and characterization of anisotropic colloidal particles, ranging from complex shapes, e.g., ellipsoids [1, 2], clusters [3], and propellers [4], to anisotropic chemical surface coatings. An important example for the latter are so-called Janus spheres [5, 6, 7, 8, 9, 10], which exhibit autonomous self-propulsion under suitable conditions and often serve as colloidal models for microswimmers. The use of such particles in experimental model systems has opened new avenues and has boosted research in the field of active matter (see, e.g., Refs. [11, 12, 13]). These successes have stimulated a large number of theoretical investigations, which have brought forth elaborate descriptions of the self-propelled motion of a single active colloid (see, e.g., Refs. [14, 15, 16, 17, 18, 19]). Depending on the question at hand, hydrodynamic effects due to solvent flows can be crucial [20, 21, 22, 23] or may be negligible. In the latter case, the active Brownian particle (ABP) model is widely used to describe the motion of active colloids [24, 17, 25]; it combines the Brownian motion of a colloidal Janus sphere with a constant propulsion velocity which randomly changes its direction. The propulsion direction is described by a unit vector $\boldsymbol{u}$ which itself performs a Brownian motion on the unit sphere. Thereby, the stochastic motion of $\boldsymbol{u}$ may evolve freely [5, 26, 27] or under the influence of gravity [28, 29, 30, 31, 32, 33] and it becomes particularly challenging in the case of shape anisotropy [34, *Doi:2005, 36, 37, 38, 39, 40, 41, *Leitmann:2017, 43]. The dynamics of active colloids becomes complex in the presence of rotation–translation coupling, e.g., due to short-time hydrodynamic friction [34, *Doi:2005, 36, 37] or due to confinement by a harmonic potential well [44] or a substrate potential [27, 45]; in such situations, the description of the rotational dynamics is crucial for predictions on the translational motion.
@@
Models for the rotational motion of molecules have traditionally been studied in the context of dynamic light scattering and dielectric spectroscopy [46, 47]. These models are typically based on the Fokker–Planck equation for orientational diffusion [48, 49, 50], which allows one to calculate correlation functions and, more generally, the statistics of the trajectories [40, 41, *Leitmann:2017, 43]. Alternatively, some form of Langevin equation is employed, which provides a description on the level of stochastic trajectories and which forms the basis of stochastic simulations. The underdamped Langevin equation is used when inertia effects are relevant [46] and, mathematically, there are no conceptual questions with this equation apart from being demanding to work with it analytically. In the context of colloidal motion, however, one would like to model the stochastic dynamics as (completely) overdamped, which requires a suitable Wiener process for the particle orientation. The corresponding Langevin equation contains a noise term that has a “multiplicative” structure [46, 34, *Doi:2005, 36, 37, 29, 31, 25, 41, *Leitmann:2017, 43, 27, 45] [see below, Eqs.˜7 and 9], which brings up questions about the underlying stochastic calculus. In particular, in its Itō form, the overdamped Langevin equation [Eq.˜9] contains a drift-like term, which may appear counter-intuitive at first sight. Another example for this mathematically well-understood situation [51, 52] is given by translational diffusion near a wall, characterized by a position-dependent diffusivity, which has lead to a re-arising apparent controversy over decades [53, 54, 55, 56, 57, 58, 59, 60, 61, 62]. The issue of multiplicative noise in rotational diffusion is entirely circumvented by modeling the particle orientation as confined to a plane, which allows one to parametrise the two-dimensional rotations by a single angle. Whereas this choice is widely adopted in models of active matter, it often appears to be a choice of convenience for simplicity rather than being rooted in physics. Notably, the linear transport behavior can sensitively depend on the dimensionality of rotational diffusion [45].
@@
In this note, we aim to clarify some mathematical aspects of rotational Brownian motion, in particular, the different forms of the overdamped Langevin equation and how they depend on the adopted definition of the stochastic integral (Itō or Stratonovich). The equations are derived from a geometric construction of Brownian motion on the unit sphere, given as a sequence of infinitesimal random rotations. The construction suggests a geometric integration scheme for rotational Brownian motion, which (i) exactly preserves the normalization constraint of the orientation vector and (ii) can be extended such that it is weakly exact for arbitrarily large time steps, i.e., it correctly generates the complete probability distributions, including the propagator [51, 63]. (A strongly convergent scheme would require that the numerical solutions also agree at the level of the individual random trajectories.)
In order to focus on the essence of the problem, we shall mainly discuss chemically anisotropic spheres (e.g., Janus spheres), thereby avoiding additional difficulties that arise from the incessant transformation of the diffusivity tensor due to the rotational motion. We then show that the results apply also to axisymmetric particles, which requires some mostly technical modifications of the derivations. In mathematical texts, such diffusion processes with constant coefficients are referred to as homogeneous in time.
@@
Finally, we mention that the literature describes a number of numerical schemes for Brownian motion of rigid bodies, which are often justified only heuristically; this includes schemes using finite rotations [64, 65], adding a constraint force [66, 67, 68, 69], discretizing the Stratonovich–Langevin equation [34, 70], or employing a symplectic splitting scheme of the underdamped Langevin dynamics [71]. Also, the algorithm presented below (Section˜IV.1) has already been used in previous work of both authors [41, *Leitmann:2017, 27, 45], without further explanation so far.
@@ II.1 Langevin equation of rotational Brownian motion with inertia
If the particle is immersed in a fluid medium, it experiences a torque $\boldsymbol{\zeta}_{R}(t)\,\boldsymbol{\omega}(t)$ due to viscous (Stokes) friction. In general, the rotational friction coefficient $\boldsymbol{\zeta}_{R}(t)$ is a tensor which is constant in a body-fixed frame [72], but is time-dependent in the space-fixed frame used here. Following Langevin’s and Ornstein’s approach to Brownian motion [73], the incessant collisions of the particle with solvent molecules are further subsumed in a random torque $\boldsymbol{\xi}(t)$ acting on the particle. Taking the collisions to be uncorrelated in a first approximation suggests to model $\boldsymbol{\xi}(t)$ as a Gaussian white noise process, which is independent of $\boldsymbol{\omega}(s)$ for $t>s$, has mean zero, and its covariance matrix is determined by the fluctuation–dissipation relation:
@@
where $k_{\text{B}}T$ denotes the thermal energy scale and $\langle\,\cdot\,\rangle$ is an ensemble average over the noise $\boldsymbol{\xi}(t)$. Foreseeing the possibility of an externally imposed torque $\boldsymbol{T}_{\text{ext}}(t)$, the total torque $\boldsymbol{T}$ has two deterministic and one random contributions so that Eq.˜1 is turned into the Euler–Langevin equation for the evolution of $\boldsymbol{\omega}(t)$:
This is a stochastic differential equation (SDE) in $\boldsymbol{\omega}(t)$ and since the white noise $\boldsymbol{\xi}(t)$ enters “additively” (i.e., it is not multiplied by a $\boldsymbol{\omega}(t)$-dependent function), there is no ambiguity about the choice of the stochastic calculus: any (consistent) definition of the stochastic integral yields the same solution [52, 51].
The above assumptions on the random torque are very well fulfilled in many use cases. However, we remind the reader that high-resolution measurements on colloidal particles have revealed deviations from Eq.˜3 in the form of persistent correlations, due to hydrodynamic memory [74, 75]. Also at short lag times, time-reversibility and smoothness of the trajectories, at the scale of molecules and atoms, implies a vanishing high-frequency friction [76] and thus $\langle|\boldsymbol{\xi}(0)|^{2}\rangle<\infty$, different from Eq.˜3. A predictive theory of the emergence of friction from Hamiltonian dynamics remains as an open problem in statistical physics. Nevertheless, one can phenomenologically account for both aspects in the framework of the generalized Langevin equation, developed in the 1960’s by Zwanzig [77], Mori [78] and others, see Ref. 79 for a recent review. Integrating out the solvent degrees of freedom in rotational Brownian motion is a program that, to our knowledge, still needs to be carried out. The use of a differentiable noise $\boldsymbol{\xi}(t)$ in Eq.˜4 would eliminate any ambiguity in the stochastic integral [80], at the price of giving up the simplicity of a Markovian dynamics in Langevin’s approach.
In Section˜V, we extend the results to axisymmetric particles, for which $\boldsymbol{u}(t)$ is naturally chosen to align with the symmetry axis of the particle. This axis fully determines the particle’s orientation such that the friction tensor $\boldsymbol{\zeta}_{R}(t)$ and moment of inertia $\boldsymbol{I}(t)$ are functions of $\boldsymbol{u}(t)$. For particles of arbitrary shape, the orientation may be tracked by following the motion of a body-fixed trihedron [34, *Doi:2005, 36], an elaboration on this generalization is left for future research.
II.2 Heuristic derivation of the overdamped limit
For small colloidal (passive or active) particles, inertia is typically negligibly small [81]. An exception are effects due to hydrodynamic memory, the latter manifesting themselves in algebraically decaying long-time tails of correlation functions [82, 74, 75, 83, 84]; however, effects due to such persistent memory are not included in the Langevin description of Brownian motion. In the absence of an external torque, one sees from Eq.˜4 that the characteristic relaxation time of the angular velocity $\omega$ is given by $\tau_{1}=I/\zeta_{R}$. For a micron-sized sphere ($a=${10}^{-3}\text{\,}\mathrm{m}\mathrm{m}$\$) in water ($\eta/\varrho=$1\text{\,}\mathrm{m}\mathrm{m}^{2}\mathrm{/}\mathrm{s}$\$), it holds $\tau_{1}=\varrho a^{2}/(15\eta)\approx${10}^{-7}\text{\,}\mathrm{s}$\$. For times $t\gg\tau_{1}$, it is thus justified to neglect the inertial term $I\dot{\boldsymbol{\omega}}$ in Eq.˜4 relative to the dissipative term $\zeta_{R}\boldsymbol{\omega}$, yielding
This simple reasoning is possible since Eq.˜4 is of the Ornstein–Uhlenbeck type, in particular, since it is linear in $\boldsymbol{\omega}(t)$ and the noise $\boldsymbol{\xi}(t)$ enters additively. The same result would be obtained from the formal solution for $\boldsymbol{\omega}(t)$, given in integral form [52], and taking the limit $\tau_{1}\to 0$.
Equation˜5 states that the angular velocity $\boldsymbol{\omega}(t)$ is statistically equivalent to a Gaussian white noise process with a deterministic bias $\boldsymbol{T}_{\text{ext}}(t)$. More precisely, $\boldsymbol{\omega}(t)$ is a Gaussian process which satisfies
$\langle\delta\boldsymbol{\omega}(t^{\prime})\otimes\delta\boldsymbol{\omega}(t)\rangle$ $=2D_{\text{R}}\boldsymbol{1}\delta(t^{\prime}-t)\,$, (6)
upon introducing the unbiased part $\delta\boldsymbol{\omega}(t)=\boldsymbol{\omega}(t)-\zeta_{R}^{-1}\boldsymbol{T}_{\text{ext}}(t)$ and the usual rotational diffusion constant $D_{\text{R}}=k_{\text{B}}T\zeta_{R}^{-1}$; from here onward, $\langle\,\cdot\,\rangle$ refers to the noise average with respect to $\boldsymbol{\omega}(t)$. We note that $\boldsymbol{T}_{\text{ext}}(t)$ may depend on $\boldsymbol{u}(t)$, e.g., for a dipole-like alignment interaction between pairs of particles, which gives rise to the subtle question if (and in which sense) $\boldsymbol{\omega}(t)$ is independent of the orientation $\boldsymbol{u}(t)$.
It is now natural to interpret the kinematic equation (2) for the orientation $\boldsymbol{u}(t)$ as an SDE with a random angular velocity $\boldsymbol{\omega}(t)$. We repeat the equation here:
$\dot{\boldsymbol{u}}(t)=\boldsymbol{\omega}(t)\times\boldsymbol{u}(t)\qquad\text{(Str)}\,$, (7)
which is one form of the overdamped Langevin equation for rotational Brownian motion of a body-fixed axis $\boldsymbol{u}(t)$. However, the white noise $\boldsymbol{\omega}(t)$ multiplies the variable $\boldsymbol{u}(t)$ and the re-interpretation of Eq.˜2 is yet ill-defined unless we specify the stochastic calculus  [^1] We emphasize that we have left open this question when solving Eq. 4., i.e., unless we give a meaning to the time integral $\int_{0}^{t}\boldsymbol{\omega}(s)\times\boldsymbol{u}(s)\,\mathrm{d}s$. This is a modeling decision that needs problem-specific insight. Recalling that $\boldsymbol{u}(t)$ is a unit vector, Eq.˜7 describes a diffusion process on a constraining manifold, and the mathematically natural choice in this situation is to interpret the time integral as a (Fisk–)Stratonovich integral (see chap. 35 in ref. 80). After having fixed the stochastic integral, the physical soundness of the model must be verified, e.g., by testing detailed balance in equilibrium (see below, Section˜VI.2).
In the physics literature, the Stratonovich interpretation of Eq.˜7 is often (tacitly and, in this case, correctly) assumed. In the case of translational motion, rigorous arguments were established for the adiabatic elimination of linear momentum from underdamped Langevin equations [53, 54], which can also be performed systematically using the the Fokker–Planck equation of the underdamped problem (Kramers equation) [86, 87]. In principle, these procedures can be adapted to the present underdamped problem in $(\boldsymbol{u},\boldsymbol{\omega})$ [Eqs.˜2 and 4] to eliminate the fast variable $\boldsymbol{\omega}$. Although both routes have recently been followed for two-dimensional rotational motion of ABPs [88], they appear to be formidable tasks in three dimensions.
For the numerical integration of Eq.˜7, a direct application of the widespread Euler–Maruyama algorithm, which is a simple and robust algorithm for Itō SDEs, breaks the conservation of the norm $|\boldsymbol{u}(t)|$. This can be fixed a posteriori by rescaling $\boldsymbol{u}(t)$ after every integration step [67] [^2] The algorithm due to Briels et al. [67] is currently implemented in the molecular dynamics simulation software LAMMPS (since release from 29 September 2021), see the command “fix Brownian/sphere” at https://docs.lammps.org/fix_brownian.html. , but such an approach appears conceptually unsatisfying. Instead, one should use a suitable discretization scheme for the Stratonovich integral, e.g., the stochastic Milstein algorithm; yet, the correct implementation of such schemes is more demanding [63, 69]. An alternative is to mix different stochastic calculi: one may cast the Stratonovich form (7) into an equivalent form that invokes the Itō integral and which is thus suitable for the Euler–Maruyama scheme [86, 46, 52]; this form includes a noise-induced term appearing as a drift [see below, Eq.˜9]. We note that all of these integration schemes preserve the normalization of $\boldsymbol{u}(t)$ only asymptotically for a vanishing integration time step, $\Delta t\to 0$.
II.3 Two-dimensional case.
For motion in two dimensional space, which is widely considered in the theoretical literature on active particles, the issue of the multiplicative noise can be circumvented since the angular velocity has only a single component. Writing $\boldsymbol{\omega}=(0,0,\omega)$, the component $u_{z}$ of $\boldsymbol{u}=(u_{x},u_{y},u_{z})$ remains unchanged under the dynamics so that we can set $u_{z}=0$ and parametrise the remaining two components $(u_{x},u_{y})=(\cos(\varphi),\sin(\varphi))$ in terms of the angle $\varphi\in[-\pi,\pi)$. The evolution of $\varphi(t)$ follows from Eq.˜2, which reads for the first vector component: $\dot{u}_{x}(t)=-\omega\,u_{y}(t).$ Combining with the ordinary chain rule, $\dot{u}_{x}(t)=-\sin(\varphi(t))\,\dot{\varphi}(t),$ yields:
$\dot{\varphi}(t)=\omega(t)\,.$ (8)
Passing to the overdamped limit, we replace $\omega(t)$ with the white noise given in Eq.˜5, which enters additively. In fact, Eq.˜8 is the simplest form of an SDE, which has an unambiguous solution with any stochastic calculus. We conclude that $\varphi(t)$ is a Brownian motion on the one-dimensional torus $\pi,\pi)$, i.e., $\boldsymbol{u}(t)$ performs a Brownian motion on the unit circle in the $xy$-plane. The same result would be obtained when starting from Eq.˜7 and observing that the classical (Leibniz) chain rule still applies for the Stratonovich calculus.
II.4 Itō form of the overdamped equation
For analytical and numerical work the use of Itō’s calculus has certain advantages and, to this end, Eq.˜7 needs to be cast into its Itō form, which reads: